OpenStudy (anonymous):

Find the solution to the equation 64(2 + x) = 44x. Using complete sentences, explain the procedure used to solve this equation

OpenStudy (anonymous):

?????

OpenStudy (tkhunny):

I was walking through the forest, one day. I stumbled across an equation, sitting under a large oak leaf. 64(2 + x) = 44x The strange equation seemed to want to be solved. I thought, "Your constants are both divisible by four.". Let's divide by 4. 16(2 + x) = 11x While stretching out on the ground near the equation, but a small brook, I contemplated additional common factors, but found none. A very busy squirrel was scooting up and down one tree, and distributing food for the winter all about the forest floor. "Wait", I thought. "Distribute"!! That's what we'll do next. 32 + 16x = 11x The ant crawling up my leg distracted me for a moment. I was a little angry with the ant. I disrupted the ants' trail so they wandered off in two directions. After a short while, as the sun was sinking low, I started to feel bad for the ants. I thought I should get their trails back together. Again, after ignoring the equation, it hit me. My variable terms were not together! Well, let's get them on the same side of that equal sign. 32 + 16x - 16x = 11x - 16x 32 = -5x Being a tough problem, I began to lose interest. "So near the end!" I thought. As is too often the case, I was a little drowsy. As the setting sun streamed through the leaves, the sunlight seemed to be divided by the leaves into glorious orange shafts of light. Light bulb!! I need to divide! That x needs closure - a feeling of completion. 32/(-5) = x I though I should check my answer, but the night moths were out and their dim patterns more than just a little interesting. I decided to let the interested reader do the checking.

OpenStudy (anonymous):

So is 32/(-5)=x the answer and how did you get there

OpenStudy (tkhunny):

Are you serious? Those were complete sentences. Did you look at the equations? Look harder!

OpenStudy (anonymous):

I read the whole thing, okay would it be -6.2

OpenStudy (tkhunny):

There should not be any more questions. I used complete sentences and solve the equation. Did you follow it? Where did you wander off if you are still lacking confidence.

OpenStudy (anonymous):

I am wondering if i would actually divide the numbers or leave it as a fraction.

OpenStudy (anonymous):

The questions is actually 64^(2+x)=46^(4x)

OpenStudy (anonymous):

not 46 just 4

OpenStudy (anonymous):

64^(2+x)=4^(4x)

OpenStudy (tkhunny):

Wow. It so helps the start with the right problem. 1) No one should care if you produce a decimal value of leave a fraction. They are the same. 2) Notice that 4^3 = 64 Then we have: 64^(2+x)=4^(4x) (4^3)^(2+x)=4^(4x) 4^[3(2+x)]=4^(4x) 4^(6 + 3x) = 4^(4x) This is a big step. Follow closely. 6 + 3x = 4x 6 = x Done. 3) You'll have to invent your own story with a squirrel in it.